Circular No. 5886

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

X-RAY NOVA IN VELA K. Borozdin, R. Sunyaev and N. Alexandrovich, on behalf of the Mir-

Kvant-TTM team, report: “The x-ray nova GRS 1009–45 (IAUC 5864, 5878) continues to be bright. The flux on Oct. 27 within the 2-6 keV spectral band was 530 mCrab (a 25 percent decrease in four weeks). The spectrum is soft and might be approximated by a power law with photon spectral index –5.0 or black body spectrum with kT = 0.56 keV. Owing to the presence of Vela X-1 in the same field TTM was able to localize the source to within a narrow strip 1′.2 wide and 1^{o}.3 long with corners at α = 10^{h}10^{m}.8, δ = −45^{o}43′.2; 10^{h}10^{m}.8, −45^{o}42′.7; 10^{h}12^{m}.4, −44^{o}26′.0 and 10^{h}12^{m}.5, −44^{o}26′.1 (equinox 1950.0). By combining the Granat/WATCH and BATSE data it is possible to reduce the strip’s length by up to 0^{o}.5.”

COMET MUELLER (1993p) Ephemeris continuation from the orbital elements on MPC 22663:

1993/94 α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Oct. 10 0^{h}15.^{m}52 +43^{o}10.^{′}8 1.844 2.702 142.^{o}0 13.^{o}1 12.1
20 23 51.46 +38 13.3 1.716 2.583 143.1 13.4 11.8
30 23 30.23 +31 56.5 1.635 2.462 137.8 15.7 11.5
Nov. 9 23 13.40 +24 50.8 1.602 2.341 127.5 19.6 11.2
19 23 01.51 +17 36.1 1.613 2.219 114.9 23.8 11.0
29 22 54.32 +10 47.1 1.658 2.096 101.8 27.4 10.8
Dec. 9 22 51.22 + 4 42.7 1.726 1.972  89.0 30.0 10.6
19 22 51.53 − 0 32.3 1.803 1.848  76.9 31.2 10.4
29 22 54.64 − 5 02.7 1.880 1.725  65.7 31.3 10.2
Jan. 8 23 00.02 − 8 57.3 1.946 1.603  55.2 30.3 10.0
18 23 07.33 −12 26.1 1.994 1.484  45.6 28.3  9.7
28 23 16.29 −15 38.7 2.018 1.368  37.1 25.7  9.4

PERIODIC COMET SCHWASSMANN-WACHMANN 1 Total CCD magnitude estimates: Oct. 11.78 UT, 16.9 (A. Nakamura,

Kuma, Japan, 0.60-m reflector); 19.14, 15.6 (P. Pravec, Ondřejov Observatory, 0.18-m reflector); 22.77, 15.3 (Nakamura); 23.68, 15.5 (T. Kojima, Chiyoda, Japan, 0.25-m Wright-Schmidt); 27.79, 13.3 (Nakamura); 31.63, 13.1 (Kojima); Nov. 1.61, 13.3 (Kojima).

1993 November 2 Brian G. Marsden